OpenStudy (anonymous):

Convert each rate using dimensional analysis. Round to the nearest tenth, if necessary. 55 mi / h = ____ ft / s a. 8.7 c. 80.7 b. 26.9 d. 806.7

OpenStudy (jhannybean):

how many minutes makeup 1 hour?

OpenStudy (anonymous):

60 minutes/hr. @Jhannybean

OpenStudy (jhannybean):

\[\cfrac{55 \ mi}{1 \ hr} \times \cfrac{ 1 \ hr}{60 \ min}\] So far. and we need to get to seconds.

OpenStudy (jhannybean):

\[\cfrac{55 \ mi}{1 \ \cancel{ hr}} \times \cfrac{ 1 \ \cancel{hr}}{60 \ min}\]

OpenStudy (jhannybean):

Now how many seconds are in 1minute?

OpenStudy (anonymous):

Okay. Do we multiply those?

OpenStudy (anonymous):

60 in 1

OpenStudy (jhannybean):

We are not done yet :P Just hang onto those numbers as they are for the time being.

OpenStudy (jhannybean):

Alrighty :D \[\large \cfrac{55 \ mi}{1 \ \cancel{ hr}} \times \cfrac{ 1 \ \cancel{hr}}{60 \ \cancel{min}} \times \cfrac{1 \ \cancel{min}}{60 \ sec}\]

OpenStudy (jhannybean):

Now we've got to convert our miles to feet... so how many feet are in 1 mile?

OpenStudy (anonymous):

5280?

OpenStudy (jhannybean):

Good job.

OpenStudy (jhannybean):

\[\large \cfrac{55 \ \cancel{mi}}{1 \ \cancel{ hr}} \times \cfrac{ 1 \ \cancel{hr}}{60 \ \cancel{min}} \times \cfrac{1 \ \cancel{min}}{60 \ sec} \times \cfrac{5280 \ ft}{1 \ \cancel{mi}}\]

OpenStudy (jhannybean):

Are you with me so far?

OpenStudy (anonymous):

Lol. I think so.

OpenStudy (jhannybean):

We have to logically place our conversions in the order of cancellation,much like we do with fractions, we want to cancel out the units to get to the conversion we are looking for.

OpenStudy (anonymous):

Okay..Lol. I'm following..

OpenStudy (jhannybean):

Therefore, we have \[\large \cfrac{55(5280) \ ft}{60 \ sec}\]

OpenStudy (anonymous):

4840

OpenStudy (jhannybean):

Yep, but it doesn't seem like one of your answer choices,so lets go back and check if we missed anything.

OpenStudy (anonymous):

k.

OpenStudy (jhannybean):

Oh sorry, it's not divided by 60, it's divided by 3600.

OpenStudy (jhannybean):

[55(5280)]/3600

OpenStudy (anonymous):

80.6?

OpenStudy (jhannybean):

it should say 80.67,not 806.7 .

OpenStudy (anonymous):

Lol. 80.7

OpenStudy (anonymous):

right? 80.7

OpenStudy (jhannybean):

Yep.

OpenStudy (anonymous):

Thank you. (:

OpenStudy (jhannybean):

No problemo.